# Generalist World Model Pre-Training for Efficient Reinforcement Learning

### v1

@@ 3.3 RL-based Fine-tuning
In our fine-tuning stage, we allow the agent to interact with the environment to collect new data $\tau_{\text{on}}^{i}=\{o^{i}_{t},a^{i}_{t},r^{i}_{t}\}_{t=0}^{T}$. The newly collected data is used to learn a reward function via supervised learning whilst also fine-tuning the world model with Eq. 1. For simplicity, we represent the concatenate of $h_{t}$ and $z_{t}$ as $s_{t}=[h_{t},z_{t}]$. The actor and critic are trained with imagined trajectories $\tilde{\tau}(s,a)$ generated by rolling out the policy $\pi_{\phi}(a|s)$ in the world model $p_{\theta}$ starting from the initial state distribution $p_{0}(s)$, which are samples from replay buffer. Following DreamerV3, the critic $v_{\phi}(V_{t}^{\lambda}\mid s_{t})$ learns to approximate the distribution
over the $\lambda$-return $V_{t}^{\lambda}$ calculated as,
$\underbrace{V^{\lambda}_{t}}_{\lambda-\text{return}}=\hat{r}_{t}+\gamma~{}\begin{cases}(1-\lambda)v_{t+1}^{\lambda}+\lambda V^{\lambda}_{t+1}&\text{if}~{}~{}t<H\\ v_{H}^{\lambda}&\text{if}~{}~{}t=H\end{cases},$ (2)
@@
where $v_{t}^{\lambda}=\mathbb{E}[v_{\phi}(V_{t}^{\lambda}\mid s_{t})]$ denotes the expected value of the distribution predicted by the critic. It is trained using maximum likelihood by minimizing
$\mathcal{L}(v_{\phi})=\mathbb{E}_{p_{\theta},\pi_{\phi}}\left[-\sum_{t=1}^{H-1}\mathrm{ln}\ v_{\phi}(V_{t}^{\lambda}\mid s_{t})\right].$ (3)
The actor $\pi_{\phi}(a_{t}|s_{t})$ is then updated by maximizing the $\lambda$-return regularized by actor entropy $\mathbf{H}[a_{t}|s_{t}]$:
@@
$\mathcal{L}(\pi_{\phi})$ $=\mathbb{E}_{p_{\theta},\pi_{\phi}}\left[\sum^{H-1}_{t=1}\left(-v_{t}^{\lambda}-\eta\cdot\mathbf{H}[a_{t}|s_{t}]\right)\right].$ (4)
In practice, we found that using the pre-trained world model alone fails to work well in many cases, especially for hard-exploration tasks. To identify the reasons, we need to rethink the role of each component in policy updates. Given the policy update rule in Eq. 4, the policy update is infected by the distribution of imagined trajectories $\tilde{\tau}(s,a)=p_{0}(s)\prod_{t=0}^{H-1}\pi_{\phi}(a_{t}|s_{t})p_{\theta}(s_{t+1}|s_{t},a_{t})$ and the reward model $r_{\theta}(s_{t},a_{t})$, i.e., the initial state distribution $p_{0}(s)$, the policy $\pi_{\phi}(a|s)$, the dynamics model $p_{\theta}(s_{t+1}|s_{t},a_{t})$, and the reward model $r_{\theta}(s_{t},a_{t})$. Intuitively, as shown in Fig. 3, the policy will learn a good policy if the imagined trajectories $\tilde{\tau}$ contain “promising” data for policy learning and the reward function $r_{\theta}(s_{t},a_{t})$ can correctly label those trajectories. The world model pre-training gives a better initialization of $p_{\theta}(s_{t+1}|s_{t},a_{t})$. However, when the online data distribution is narrow – as in hard exploration tasks – the initial state distribution $p_{0}(s)$ can also be narrow, which prevents the actor from reaching high-reward states during imagination. Also, the reward function $r_{\theta}(s,a)$ trained on the narrow online data distribution will incorrectly assign rewards for transitions generated by the world model pre-trained on a broader distribution. That is, the agent fails to improve the policy even if “promising” trajectories exist in $\tilde{\tau}$, since incorrect rewards are assigned. Furthermore, fine-tuning the world model $p_{\theta}(s_{t+1}|s_{t},a_{t})$ on the narrow distribution causes catastrophic forgetting. In the following, we improve the agents’ performance by reusing non-curated offline data for i) augmenting the initial state distribution $p_{0}(s)$, ii) collecting diverse online trajectories, and iii) improving the reward predictions. See Fig. 3 for an illustrative motivation.
Retrieval-based Experience Rehearsal
Figure 3: Motivation of using experience rehearsal and execution guidance. During fine-tuning, experience rehearsal augments the initial state $p_{0}(s)$ used for model rollout, which enables the model rollout to start from informative states. Execution guidance helps the agent to collect data that is close to the offline data distribution, which enriches the diversity of the collected data and improves the reward model’s accuracy on a broader state distribution.
Experience replay is a simple yet effective approach to prevent catastrophic forgetting (Khetarpal et al., 2022) as well as improving policy learning (Ball et al., 2023; Li et al., 2023). In the RL setting, Ball et al. (2023) demonstrates promising results by replaying offline data to improve sample efficiency but requires the offline data to be reward-labeled. ExPLORe (Li et al., 2023) reuses single-embodiment but reward-free offline data to encourage exploration by labeling offline data with UCB rewards. Unlike the replay scheme used in ExPLORe, where the offline datasets are relatively small and well-structured, directly replaying the non-curated offline data can be challenging. Our dataset can be 100$\sim$ 100$\times$ 100 × larger and consist of different tasks and embodiments. As such, directly labeling the offline dataset with UCB reward or even replaying the dataset can be infeasible.
To enable the usage of the non-curated offline data, we propose a simple retrieval-based approach to filter the trajectories that are close to the downstream tasks. Specifically, we retrieve a subset of trajectories from the non-curated offline data based on the neural feature distance between the online samples and the trajectories in the offline dataset. We calculate the distance as :
$\mathbf{D}=||\text{e}_{\theta}(o_{\text{on}})-\text{e}_{\theta}(o_{\text{off}})||_{2},$ (5)
where $\text{e}_{\theta}$ is a feature extractor and $o_{\text{on}}$ and $o_{\text{off}}$ are images from the initial observation of trajectories from the online buffer and the offline dataset, respectively. In practice, we found that neural representation matters for retrieval and the encoder learned via world model pre-training works better than the general visual model R3M (Nair et al., 2022).
To efficiently search over the whole dataset, we pre-build the key-value pairs between the trajectories ID and its neural features and use Faiss (Douze et al., 2024) for similarity search. Once the key-value database is built, the buffer retrieval procedure only takes seconds. The retrieved offline buffer, although reward-free, contains embodiment-specific data and is feasible to use. The retrieved offline data can be used to prevent catastrophic forgetting in the world model and to augment the initial state distribution $p_{0}(s)$ used for trajectory imagination.
Figure 4: WPT enables sample-efficient RL training. We include 6 tasks from DMControl and Meta-World with pixel inputs. The mean and the corresponding 95% confidence interval across 3 seeds are plotted. Full results on 72 tasks are included in App. D.
@@
The common practice in RL training is to initialize the replay buffer with a random policy and gradually collect new data by interacting with the environment using the training policy. When using a pre-trained world model, it is preferred to start the training from a distribution that is close to the offline data distribution. This has three benefits: First, offline data usually contains useful information for policy training, such as near-expert trajectories or a broader state-action distribution. Second, the distribution shift between the offline and online data may ruin the pre-trained weights, so bringing the initial distribution close to the offline data is a good strategy to prevent this. Thirdly, the reward function can get better predictions for states that are close to the offline data distribution. This improves policy learning as the initial states used to generate the imagined trajectories $\tilde{\tau}$ are augmented with offline samples.
To guide the online data collection such that it remains close to the offline data distribution, we simply train a prior actor $\pi_{\text{bc}}$ via behavior cloning on the retrieved buffer. During online data collection, we switch between the policy prior $\pi_{\text{bc}}$ and the RL agent $\pi_{\phi}$ according to a pre-defined schedule. Specifically, at the beginning of an episode, we decide whether to use $\pi_{\text{bc}}$ according to the scheduled probability. If the $\pi_{\text{bc}}$ is selected to use, we then randomly select the starting time step $t_{\text{bc}}$ and duration $H$ of using $\pi_{\text{bc}}$. $\pi_{\phi}$ is then used for the rest of the time steps. This procedure is similar to JSRL (Uchendu et al., 2023) but with a couple of differences: i) WPT leverages non-curated offline data while JSRL uses task-specific offline data and ii) WPT shows the superiority of a model-based approach for leveraging offline data while JSRL is a model-free approach. As a result, we show that WPT obtains better performance than JSRL on tested tasks. The full algorithm is presented in App. C.
@@ 4 Experiments
In this section, we investigate our WPT in two different RL settings. In the first setting, we focus on how the proposed WPT helps to improve the sample efficiency of RL training. In the second setting, we investigate whether WPT enables fast task adaptation, which is particularly important to empower the agent’s life-long learning ability. Lastly, we conduct an ablation study to investigate the role of proposed design choices.
4.1 Dataset
Our dataset consists of data from two different benchmarks: DeepMind Control Suite (DMControl) and Meta-World, visualized in Fig. 2 and Fig. 9 for the full list. For DMControl, we include 10k trajectories covering 5 embodiments collected by unsupervised RL agents (Pathak et al., 2017; Rajeswar et al., 2023), which are trained via curiosity without any task-related information. For Meta-World, we collect 50k trajectories spinning 50 tasks by executing the pre-trained RL agents from TDMPCv2 (Hansen et al., 2024) with injected Gaussian noise with $\sigma=[0.1,0.3,0.5,1.0,2.0]$ to mimic imperfect demonstrations. In total, the offline dataset consists of 60k trajectories with 10M state-action pairs from 6 different embodiments.
4.2 WPT Enables Sample-efficient Fine-tuning
@@
Tasks   We evaluate our method on six selected pixel-based continuous control tasks from DMControl and Meta-World. The chosen tasks cover different challenges in RL. Except for learning from high-dimensional observations, we include a hard exploration version of Cheetah Run by setting zero rewards below a reward threshold and applying an action penalty. We include Walker Run and Quadruped Walk for complex dynamics. We further include challenging manipulation tasks from MetaWorld.
Baselines   In our experiments, we consider four different baselines for leveraging reward-free offline data. Note that as these baselines are not designed to handle multi-embodiment offline data, we preprocess the offline data to only contain task-related data for the baselines. Although this is an unfair comparison for WPT as baselines eliminate the difficulty raised with the large non-curated dataset, we still outperform these baselines by a large margin on tested tasks. The compared baselines are:
  • R3M pre-trains visual representation aiming to boost policy learning. We use a ResNet-18 backbone and train additional policy and value heads via RL. We include this baseline to compare with a representative approach for visual pre-training.
  • UDS suggests to label offline data with zero. UDS is an offline RL method, which cannot be directly used for comparison. We instead set zero reward for offline data and use RLPD (Ball et al., 2023) for policy training.
  • ExPLORe labels offline data with UCB reward. We follow the implementation from Li et al. (2023) but use ensembles for reward estimation, which has better performance.
  • JSRL-BC (Uchendu et al., 2023) collects online data with both the training policy and a prior policy trained on offline data via IL. We include JSRL-BC to compare with different ways of using offline data.
@@
Figure 5: WPT enables fast task adaptation. We train an RL agent to control an Ant robot from DMControl to finish a series of tasks incrementally. WPT outperforms the widely used baseline by a decent margin by properly leveraging non-curated offline data.

### v2

@@ 3.3 RL-based Fine-Tuning with Rehearsal and Guidance
In our fine-tuning stage, the agent interacts with the environment to collect new data $\tau_{\text{on}}^{i}=\{o^{i}_{t},a^{i}_{t},r^{i}_{t}\}_{t=0}^{T}$. This data is used to learn a reward function via supervised learning while fine-tuning the world model with Eq. 1. For simplicity, we represent the concatenation of $h_{t}$ and $z_{t}$ as $s_{t}=[h_{t},z_{t}]$. The actor and critic are trained using imagined trajectories $\tilde{\tau}$ generated by rolling out the policy $\pi_{\phi}(a\,|\,s)$ in the world model $p_{\theta}$, starting from initial states $p_{0}(s)$ sampled from the replay buffer. The critic $v_{\phi}(V_{t}^{\lambda}\,|\,s_{t})$ learns to approximate the distribution over the $\lambda$-return $V_{t}^{\lambda}$, calculated as:
$\underbrace{V^{\lambda}_{t}}_{\lambda-\text{return}}=\hat{r}_{t}+\gamma~{}\begin{cases}(1-\lambda)v_{t+1}^{\lambda}+\lambda V^{\lambda}_{t+1}&\text{if}~{}~{}t<H\\ v_{H}^{\lambda}&\text{if}~{}~{}t=H\end{cases}$ (2)
@@
where $v_{t}^{\lambda}=\mathbb{E}[v_{\phi}(V_{t}^{\lambda}\,|\,s_{t})]$ denotes the critic’s expected value predicted. The value function $v_{\phi}$ and actor $\pi_{\phi}$ are updated by maximizing the log likelihood and entropy-regularized $\lambda$-return respectively:
@@
$\mathcal{L}(v_{\phi})=\mathbb{E}_{p_{\theta},\pi_{\phi}}\left[-\sum_{t=1}^{H-1}\mathrm{ln}\ v_{\phi}(V_{t}^{\lambda}\,|\,s_{t})\right],~{}~{}~{}\mathcal{L}(\pi_{\phi})$ $=\mathbb{E}_{p_{\theta},\pi_{\phi}}\left[\sum^{H-1}_{t=1}\left(-v_{t}^{\lambda}-\eta\cdot\mathbf{H}[a_{t}\,|\,s_{t}]\right)\right].$ (3)
Figure 2: Visualization of Distribution Mismatch. Left: At the early stage of fine-tuning, there is a distribution shift between offline data used for world model pre-training and online data used for RL fine-tuning, which hurts performance. Middle: Experience rehearsal mitigates the distributional shift issue. Right: Quantitatively, at the early stage of fine-tuning, experience rehearsal reduces the Wasserstein distance between the online data and both the offline and expert data.
Motivation
While previous methods typically discard non-curated offline data during fine-tuning [26, 28, 53], we find that relying solely on a pre-trained world model often fails, particularly on hard-exploration tasks. To understand why, we analyze the Shelf Place task from Meta-World as an illustrative task by visualizing the distributions of offline data $\mathcal{D}_{\text{off}}$ used for world model pre-training and online data $\mathcal{D}_{\text{on}}$ collected during early RL training in Fig. 2. The t-SNE plot in Fig. 2 (left) reveals a distribution mismatch between $\mathcal{D}_{\text{off}}$ and $\mathcal{D}_{\text{on}}$, leading to three key issues: (i) The world model’s accuracy degrades on states visited by the early-stage policy, especially when the offline data distribution is narrow, hurting sample efficiency. (ii) For hard exploration tasks, the agent struggles to reach high-reward regions, causing the world model to be fine-tuned on a narrow online data distribution and leading to catastrophic forgetting. (iii) The policy update in Eq. 3 relies on imagined trajectories $\tilde{\tau}=p_{0}(s)\prod_{t=0}^{H-1}\pi_{\phi}(a_{t}\,|\,s_{t})p_{\theta}(s_{t+1}\,|\,s_{t},a_{t})$, where $p_{0}(s)$ is sampled from $\mathcal{D}_{\text{on}}$. A narrow $p_{0}(s)$ limits the world model to rollout promising trajectories for policy updates. To address these challenges, we introduce two key components: i) experience rehearsal, which mitigates distributional shift by retrieving task-relevant trajectories from non-curated datasets (Fig. 2 middle, right), and ii) execution guidance, which encourages exploration by steering the agent toward regions where the world model has high confidence.
Experience Rehearsal
Prior work like RLPD [35] and ExPLORe [36] has shown that replaying offline data can boost RL training. However, these methods use small, well-structured offline datasets. In our setting, directly replaying non-curated offline data is infeasible since our datasets are $\sim$100× larger and contain diverse tasks and embodiments.
We propose retrieving task-relevant trajectories $\mathcal{D}_{\text{retrieved}}=\{\tau^{i}_{\text{retrieved}}\}^{N}_{i=1}$ from the non-curated offline data based on neural feature distance between online samples and offline trajectories. This filters out irrelevant trajectories, creating a small dataset of task-relevant data. Specifically, we compute:
$\mathbf{D}=\|\text{e}_{\theta}(o_{\text{on}})-\text{e}_{\theta}(o_{\text{off}})\|_{2},$ (4)
where $\text{e}_{\theta}$ is the encoder learned during world model pre-training, and $o_{\text{on}}$ and $o_{\text{off}}$ are initial observations from trajectories in the online buffer and offline dataset, respectively. For efficient search the top-k similar trajectories, we pre-compute key-value pairs mapping trajectory IDs to neural features and use Faiss [61], enabling retrieval in seconds.
The retrieved data is replayed during fine-tuning, so-called experience rehearsal. The retrieved data serves three purposes, as shown in Fig. 1. First, it prevents catastrophic forgetting by continuing to train the world model on relevant pre-training data, particularly important for hard exploration tasks with narrow online data distributions. Second, it augments the initial state distribution $p_{0}(s)$ during model rollout, enabling the world model to rollout promising trajectories for policy learning. Third, as described below, it enables learning a policy prior for execution guidance. Unlike RLPD and ExPLORe, we do not use this data to learn a Q-function, eliminating the need for reward labels.
@@
Standard RL training initializes the replay buffer with random actions and collects new data through environment interaction using the training policy. However, offline data often contains valuable information like near-expert trajectories and diverse state-action coverage that should be utilized during fine-tuning. Additionally, distribution shift between offline and online data can degrade pre-trained model weights, making it important to guide the online data collection toward the offline distribution at the early training stage.
To achieve this, we train a prior policy $\pi_{\text{bc}}$ via behavioral cloning on the retrieved offline data $\mathcal{D}_{\text{retrieved}}$. During online data collection, we alternate between this prior policy $\pi_{\text{bc}}$ and the RL policy $\pi_{\phi}$ according to a pre-defined schedule. Specifically, at the start of each episode, we probabilistically select whether to use $\pi_{\text{bc}}$. If $\pi_{\text{bc}}$ is selected, we randomly choose a starting timestep $t_{\text{bc}}$ and duration $H$ during which $\pi_{\text{bc}}$ is active, with $\pi_{\phi}$ used for the rest timesteps.
While this approach shares similarities with JSRL [33], our method differs in three key aspects: i) we leverage non-curated rather than task-specific offline data, ii) we demonstrate the benefits of a model-based approach over JSRL’s model-free framework, and iii) we randomly switch between policies mid-episode rather than only using $\pi_{\text{bc}}$ at episode start. The complete algorithm and theoretical analysis can be found in App. H and App. B, respectively.
@@ 4 Experiments
Figure 3: Left: Quantitative comparison across 72 diverse tasks from Meta-World and DMControl. GSA achieves a 102.8% relative improvement in aggregate score over learning-from-scratch baselines when using the same sample budget (150k). It also matches the performance of baselines even when they are trained with substantially more samples (see App. I for full results). Right: Learning curves on representative challenging locomotion and robotic manipulation tasks. GSA consistently outperforms state-of-the-art methods that leverage offline data by a decent margin. We plot the mean and corresponding 95% confidence interval.
In the experiments, we aim to answer the following questions: (i) How does GSA compare to state-of-the-art methods that leverage offline data and train-from-scratch baselines in terms of sample efficiency and final performance? (ii) How does GSA compare to other leading model-based approaches that utilize offline data? (iii) How effectively does GSA adapt to new tasks in a continual learning setting? We further conduct detailed ablation studies to thoroughly evaluate our method.
Tasks
@@
We evaluate our method on pixel-based continuous control tasks from DMControl and Meta-World. The chosen tasks include both locomotion and manipulation tasks covering different challenges in RL, including high-dimensional observations, hard exploration, and complex dynamics. We use three random seeds for each task.
Dataset
Our dataset consists of data from two benchmarks: DeepMind Control Suite (DMControl) and Meta-World, visualized in Fig. 12. For DMControl, we include 10k trajectories covering 5 embodiments collected by unsupervised RL agents [28, 62], trained via curiosity without task-related information. For Meta-World, we collect 50k trajectories across 50 tasks using pre-trained RL agents from TDMPC v2 [18]. To create a diverse dataset with varying quality, we inject Gaussian noise with $\sigma\in\{0.1,0.3,0.5,1.0,2.0\}$ during execution. Combined with the DMControl data, our complete offline dataset contains 60k trajectories (10M state-action pairs) spanning 6 embodiments.
4.1 GSA Improves Sample Efficiency Across Diverse Tasks
Comparison with Methods that Leverage Offline Data
We compare GSA against several state-of-the-art methods that leverage reward-free data to improve RL training: (i) R3M [11], a visual representation pre-training approach that serves as our baseline for comparing pre-trained visual features using non-curated offline data. (ii) UDS [46], which assigns zero rewards to offline data. Since UDS is an offline RL method, we combine it with RLPD [35] for policy training. (iii) ExPLORe [36], which labels offline data using UCB rewards. We enhance the original implementation with reward ensembles. (iv) JSRL-BC [33], which collects online data using a mixture of the training policy and a behavior-cloned prior policy learned from offline data. As the compared baselines cannot handle multi-embodiment data like GSA, we preprocess the offline data to only include task-relevant trajectories for them. Despite the baselines having access to better-structured data, GSA still significantly outperforms all baselines across the tested tasks. See App. G for the details of baselines.
Results   Fig. 3 (right) shows comparison results with baselines. Our method outperforms all compared baselines by a large margin. Compared to R3M, GSA shows the importance of world model pre-training and reusing offline data during fine-tuning, versus representation learning alone. R3M fails to improve sample efficiency on most tasks, consistent with findings in Hansen et al. [63].
UDS and ExPLORe reuse offline data by labeling it with zero rewards and UCB rewards, respectively, and concatenating it with online data for off-policy updates. UDS shows only slightly better performance on Walker Run compared to R3M and JSRL-BC, demonstrating the ineffectiveness of zero-reward labeling. ExPLORe performs better on 2/3 locomotion tasks and shows progress on challenging manipulation tasks, but GSA still significantly outperforms it, demonstrating the superiority of leveraging a generalist world model and properly reusing offline data during fine-tuning.
GSA also clearly outperforms JSRL-BC. JSRL-BS’s performance heavily depends on offline data distribution. While JSRL-BC can perform well when a good prior actor can be extracted from offline data, it struggles with non-expert trajectories, showing only marginal improvements over other baselines on Cheetah Run Hard, Assembly, and Shelf Place tasks. In contrast, GSA effectively leverages non-expert offline data. For example, on Quadruped Walk, GSA benefits from exploratory offline data collected by unsupervised RL, enabling pixel-based control within just 100 trials.
Comparison with Training-from-Scratch Methods
We compare GSA with two widely used training-from-scratch baselines: DrQ v2 and Dreamer v3, representing model-free and model-based approaches, respectively.  Figure 3 (left) and App. I show comparison results on 22 locomotion and 50 robotic manipulation tasks with pixel inputs from DMControl and Meta-World benchmarks. With 150k online samples, GSA achieves a 102.8% relative improvement in aggregate score compared to DrQ v2 and Dreamer v3, matching their performance obtained with 3.3-6.7$\times$ more samples (500k for DMControl, 1M for Meta-World). For example, GSA enables an Ant robot to walk forward within 100 trials, while widely used learning-from-scratch baselines require 10-30$\times$ more samples. Furthermore, GSA achieves promising performance on hard exploration tasks where learning-from-scratch baselines fail, such as challenging Meta-World manipulation tasks and hard DMControl tasks.
Figure 4: Comparison with other world model pre-training methods. GSA outperforms state-of-the-art model-based methods without relying on techniques used in iVideoGPT, such as reward shaping and demonstration-based replay buffer initialization.
Comparison with Other Model-Based Methods
While most multi-task/multi-embodiment world models focus on visual prediction [22, 23] or imitation learning [19, 20], some works like Seo et al. [24]Wu et al. [25], and iVideoGPT [26] investigate world model pre-training with in-the-wild data for RL. These methods typically focus on designing novel or scalable model architectures to leverage the offline data, but lack mechanisms to better leverage offline data during RL fine-tuning. Furthermore, due to the cost of RL training, these methods are usually evaluated on limited task sets, making the effectiveness of the pre-trained world model unclear on diverse tasks.
Figure 4 compares our method with world model pre-training approaches. The baseline results are from the iVideoGPT paper. Despite extensive pre-training on diverse manipulation data, iVideoGPT and pre-trained DreamerV3 show only marginal improvements over training-from-scratch baselines. In contrast, GSA clearly accelerates RL training by properly leveraging non-curated offline data during both pre-training and fine-tuning. Notably, baselines in Fig. 4 use reward shaping and expert replay buffer pre-filling, while GSA uses none of these tricks yet achieves superior performance. This highlights that (i) non-curated offline data contains useful information for RL fine-tuning, and ii) GSA can effectively leverage such data. Furthermore, GSA could potentially be combined with iVideoGPT to leverage even more diverse offline data in future work.
4.2 GSA Enables Fast Task Adaptation
@@
Figure 5: GSA enables fast task adaptation. We train an RL agent to control an Ant robot from DMControl to complete a series of tasks incrementally. GSA significantly outperforms the widely used baseline PackNet by properly leveraging non-curated offline data.
We investigate GSA’s benefits for continual task adaptation, where an agent must incrementally solve a sequence of tasks. While similar to continual reinforcement learning (CRL) or life-long RL [64, 65], we use a simplified setting with a limited task set. Note that CRL has a broad scope; assumptions and experiment setups vary among methods, making it difficult to set up a fair comparison with other methods. Rather than proposing a state-of-the-art CRL method, we aim to demonstrate that GSA offers an effective approach to leverage previous data that also fits the CRL setting.
Setup & Baselines
We set our continual multi-task adaptation experiment based on the Ant robot from the DMControl. Specifically, the agent sequentially learns stand, walk, run, jump, roll, and roll fast tasks with 300K environment steps per task. To have a fair comparison, i.e., having comparable model parameters and eliminating the potential effects from pre-training on other tasks, we pre-train a small world model only on the Ant domain. During training, the agent can access all previous experiences and model weights. We compare against a widely used baseline PackNet [66], which iteratively prunes actor parameters while preserving important weights (with larger magnitude) for previous skills. For each new task, PackNet fine-tunes the actor model via iterative pruning while randomly reinitializing the critic model since rewards are not shared among tasks.
Results
Figure 5 shows GSA significantly outperforms PackNet, enabling adaptation within 100 trials per task. With limited samples, PackNet achieves only 20–60% of GSA’s episodic returns. We attribute GSA’s superior performance to its ability to leverage the diverse offline data through both world model pre-training and fine-tuning with experience rehearsal, and execution guidance.
Figure 6: Ablation study on key components. “P” represents world model pre-training, “ER” means experience rehearsal, and “G” represents execution guidance. The combination of pre-trained generalist world models with retrieval-based experience rehearsal and execution guidance boosts RL performance across diverse tasks.
4.3 Ablations
Role of Each Component
We now analyze each component’s contribution using the same set of tasks from Sec. 4.1. As shown in Fig. 6, world model pre-training shows promising results when the offline data consists of diverse trajectories, such as data collected by exploratory agents (Walker Run), while it fails to work well when the offline data distribution is relatively narrow as in the Meta-World tasks. We found that experience rehearsal and execution guidance stabilize training and improve performance on hard exploration tasks like Cheetah Run Hard and challenging manipulation tasks from Meta-World. This addresses (i) world model pre-training alone, failing to fully leverage rich state and action information from the non-curated offline data and (ii) distributional shift between offline and online data during fine-tuning hurts the learning. The proposed retrieval-based experience rehearsal and execution guidance help utilize offline data and accelerate exploration, which together enable GSA to achieve strong performance on a wide range of tasks.
Figure 7: Comparison of execution guidance versus uncertainty-based reward labeling. GSA demonstrates the effectiveness of using execution guidance over uncertainty-based reward labeling on challenging robotic manipulation tasks.
Figure 8: Impact of fine-tuning different world model components.
Comparison with Uncertainty-Aware Reward Function
To leverage reward-free offline data, ExPLORe [36] proposes to label offline data with uncertainty-based rewards. To demonstrate the effectiveness of GSA, we compare it with uncertainty-based rewards. Specifically, instead of using execution guidance, we use Optimistic Thompson Sampling (OTS) [67] to label the imagined trajectories via model rollout. As shown in Fig. 7, our method outperforms the variant using OTS on hard exploration tasks, Assembly and Stick Pull, by a large margin, showing the effectiveness of using execution guidance.
Comparison of Fine-Tuning Different Components
We now investigate the role of different components in the world model during fine-tuning. We use the Quadruped Walk task as a representative task for the investigation. As shown in Fig. 8, encoder and decoder as well as latent dynamics play compatible roles during fine-tuning. Fine-tuning the full world model yields the best performance on the tested task. The full world model is fine-tuned by default in our experiments.
5 Conclusion
We propose GSA, a simple yet efficient approach to leverage ample non-curated offline datasets consisting of reward-free, mixed-quality data collected across multiple embodiments. GSA pre-trains a generalist world model on the non-curated data and adapts to downstream tasks via RL. We show that naive fine-tuning of world models fails to accelerate RL training due to distributional shift and propose two essential techniques - experience rehearsal and execution guidance - to mitigate this issue. Equipped with these techniques, we demonstrate that generalist world models pre-trained on non-curated data are able to boost RL’s sample efficiency across a broader range of locomotion and robotic manipulation tasks. We compared GSA against a wide set of baselines, including two widely used training-from-scratch methods, five methods that utilize offline data, and one continual learning method. Our GSA consistently delivers strong performance over these baselines. Extensive ablation studies reveal the effectiveness of the proposed techniques. While promising, GSA can be improved in multiple ways: extending to real-world applications, leveraging in-the-wild offline data, and exploring novel world model architectures.
Acknowledgments
We acknowledge CSC – IT Center for Science, Finland, for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through CSC. We acknowledge the computational resources provided by the Aalto Science-IT project. We acknowledge funding from the Research Council of Finland (353138, 362407, 352788, 357301, 339730). Aidan Scannell was supported by the Research Council of Finland, Flagship program Finnish Center for Artificial Intelligence (FCAI).
Appendices
[…]
Appendix A More Results
Comparison with Imitation Learning Baseline
To demonstrate the mix-quality property of the non-curated dataset, we compare GSA with Diffusion Policy, a widely used imitation learning approach by modeling the agent with diffusion models. From Fig. 9, we can see that due to the dataset consisting of non-expert data, the diffusion policy fails to demonstrate satisfactory results, while GSA can effectively utilize the offline data.
Figure 9: Comparison with Diffusion Policy. GSA can effectively handle non-curated offline data while the imitation learning baseline fails.
Full Results of Comparison with iVideoGPT
We compare with other model-based approaches on tasks used in iVideoGPT [26]. We show that GSA outperforms the baselines without using reward shaping and pre-filling the replay buffer with demonstrations. This highlights that although non-curated, the offline data can clearly boost RL training, and GSA can effectively use the information in the data.
Figure 10: Comparsion with model-based approaches for leveraging offline data.
Role of Each Component
We offer inter-quartile mean(IQM) and optimality gap for the ablation study of the role of each proposed component. Together with the retrieval-based experience rehearsal and execution guidance, a pre-trained generalist world model boosts RL performance on a wide range of tasks.